# SepLLM: Accelerating Large Language Models by Compressing One Segment into One Separator

### v1

@@ 1 Introduction
  • We have made our implementation publicly available at sepllm.github.io. Our codebase supports efficient multi-node distributed training with accelerated attention module Sep-Attention and also includes numerous existing Fusion Operators to accelerate the training process, such as fused rope (Su et al., 2023), fused layer norm, etc.
@@ 3.3 Tailored Streaming Design
Figure 4 illustrates the SepLLM’s processing architecture for streaming inference applications. The diagram depicts multiple iterations, with each row representing a distinct processing cycle. The system simultaneously maintains four specialized cache blocks: Initial Cache, Separator Cache, Past Window Cache, and Local Window Cache. Specifically, Initial Cache captures the attention sink proposed by Xiao et al. (2024b). Local Window and Past Window Caches store the KVs for consecutive tokens, with Past Window Cache serving as an overflow buffer for the Local Window Cache. Separator Cache retains the KVs for separators which contain condensed segment information.
To describe the cache management policies, we denote the runtime usage of the four caches as $Size_{\text{init}}$, $Size_{\text{sep}}$, $Size_{\text{past\_w}}$, and $Size_{\text{local\_w}}$, respectively. The runtime usage across all KV caches is defined as $Size_{\text{run}}\coloneqq Size_{\text{init}}+Size_{\text{sep}}+Size_{\text{past\_w}}+Size_{\text{local\_w}}$, which satisfies $Size_{\text{run}}\leq\textbf{{c}}$. The number of contiguous Neighboring tokens is defined as $\textbf{{n}}\!\coloneqq\!Size_{\text{past\_w}}\!+\!Size_{\text{local\_w}}$. Notably, n is a function of the input sequence length $m$ rather than a fixed hyperparameter for streaming setting. For clarity, we detail the preset hyperparameters of this caching system as follows.
  • c: The maximum capacity of the entire KV caches.
@@
  • w: The maximum capacity of Local Window Cache. Notably, w is also the minimum value of n after the runtime KV caches usage $Size_{\text{run}}$ reaches c.
@@
$\lim\limits_{m\to\infty}\!\overline{\text{Size}_{\text{run}}}$ $=\lim\limits_{m\to\infty}\!\bar{\textit{{n}}}_{m}+\textbf{{a}}+\textbf{{s}}$
@@
Our positional encoding strategy for streaming settings is the same as the state-of-the-art StreamingLLM (Xiao et al., 2024b), designed specifically for infinite-length inputs, where we focus on positions within the caches instead of those in the original text.
@@ 4.1 Experimental Settings
We evaluate our proposed SepLLM on the following tasks, i.e., training-free, training from scratch, posting-training and streaming applications.
@@
Two popular model families, i.e., Pythia (Biderman et al., 2023) and Llama-3 (Dubey et al., 2024), are employed for evaluation. Specifically, Pythia-160m-deduped is used in the training from scratch tasks since the model, data, configurations, and checkpoints are all open-source and the training results are reproducible. As for post-training settings, we take Pythia-1.4B-deduped as our base model. Even though Llama-3 exhibits powerful performance on various downstream task, these training experimental details are not available. Therefore, we only use Llama-3 for training-free and streaming tasks.
@@
In the training from scratch and post-training tasks, the deduplicated Pile (Gao et al., 2020) is utilized for training, which contains about 207B tokens. And all other configurations are the same as the corresponding settings as Pythia (Biderman et al., 2023). Specifically, the training epoch is set to 1.5 epoch (143000 steps with the global batch size as 1024), which means about 300B tokens in total are utilized for training from scratch, which is identical to Pythia (Biderman et al., 2023).
@@
Table 1: Evaluation results and average running time KV cache usage for training-free experiments on GSM8K-CoT 8-shots and MMLU 5-shots. For SepLLM and StreamingLLM, three initial tokens’ KV is kept for this experiment. r.KV(%) here represents the ratio of KV usage at runtime for the respective method compared to Vanilla.
@@ 4.2 Training-free
The representative and commonly-used GSM8K-CoT (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021)) are adopted. GSM8K-CoT (Cobbe et al., 2021) tests a model’s ability to solve mathematical problems by evaluating its reasoning and step-by-step problem-solving skills. CoT (Wei et al., 2022) means the ability to simulate a reasoning process by breaking down complex problems into a series of logical steps. And the commonly-used 8 shots are adopted. MMLU (Hendrycks et al., 2021) assesses a model’s general knowledge and reasoning across a wide range of subjects, such as history, science, mathematics and so on. The common 5-shot setting is used for MMLU.
@@
Table 2: The performance of downstream tasks and the usage of running-time KV cache in the training from scratch setting.
@@ Results.
The experimental results for training-free are shown in Table 1. “Vanilla” represents the original Llama-3 model with full attention, while “StrmLLM” represents StreamingLLM (Xiao et al., 2024b). n means the number of KV for neighboring tokens we retain. For SepLLM, all the KV for Special Tokens are kept and for the setting SepLLM (n=256), we find that SepLLM exhibits comparable performance in both multi-step mathematical CoT task and multidisciplinary knowledge reasoning tasks, when compared to the full-attention Llama-3. SepLLM achieves this using only 47.36% of KV utilized by the original Llama-3 for reasoning, indicating SepLLM’s capability of modeling contexts requiring both multi-step logical analysis and those involving multi-domain knowledge reasoning while retaining only 50% original KV.
StrmLLM (n=256) setting corresponds to removing all separators’ KV from  SepLLM (n=256) setting, except for those in Neighboring and Initial tokens. We observe a noticeable decrease in both mathematical analysis and multidisciplinary knowledge reasoning abilities of StrmLLM (n=256). StrmLLM (n=256) utilizes only 26.00% and 37.73% of the KV for the GSM8K and MMLU tasks, respectively, which are less than  SepLLM (n=256) (47.36% and 44.61% respectively). Consequently, we increase the n of StrmLLM to 380, aligning the kept KV on GSM8K to be equal to  SepLLM (n=256) (approximately 47%, while on MMLU task, StrmLLM (n=380) retains 52.5% of the KV, significantly higher than  SepLLM (n=256)). This leads to improved performance compared to StrmLLM (n=256). However, it still remains lower than the full-attention Llama-3 and  SepLLM (n=256). This indicates that the KV of separators indeed encapsulate information contained within their respective segments, and removing them significantly impacts the Transformer’s understanding and reasoning abilities.
@@
Figure 5: Training loss curves for training from scratch. 5(b) shows the ratio of the loss values of different methods to that of Vanilla with respect to FLOPs.
@@ 4.3 Training from Scratch
We train the original Pythia-160m-deduped model as well as the Pythia-160m-deduped model modified with the SepLLM (and StreamingLLM) architecture on the Pile dataset for 143,000 steps using a global batch size of 1024 (involving approximately 300B tokens in total for training). All training configurations are consistent with Pythia (Biderman et al., 2023) (see Section 4.1). And following Pythia (Biderman et al., 2023), we conduct testing on the following downstream tasks: ARC-Challenge and ARC-Easy (Clark et al., 2018), LAMBADA (Paperno et al., 2016) (for Perplexity and Accuracy), LogiQA (Liu et al., 2021), PIQA (Bisk et al., 2020), SciQA (Welbl et al., 2017). From the loss curves depicted in Figure 5 and the downstream performance in Table 2, we draw the following analysis.
@@
We find that employing a certain hybrid architecture is beneficial for both the training loss and the performance on downstream tasks. For instance, by modifying only the first self-attention layer to full attention in the experiment corresponding to SepLLM (n=64) (denoted as SepLLM (n=64,H)), there is a moderate optimization in both the training process and downstream tasks. If both the first and last attention layers are changed to full attention (denoted as SepLLM (n=64,H/T)), this optimization becomes more pronounced. For example, LAMBADA perplexity decreases from 40.08 for SepLLM (n=64) to 36.54 for SepLLM (n=64,H) and 33.41 for SepLLM (n=64,H/T).
@@
The experiment with the setting StrmLLM (n=64) corresponds to SepLLM (n=64), but does not consider separators other than those in Neighboring and Initial tokens. We observe a significant slowdown in the training loss decrease for StrmLLM (n=64), and its performance deteriorates across various downstream tasks. This indicates that the KV corresponding to separators contain information about the segment they belong to, which is beneficial for predicting subsequent tokens.
@@
Table 3: The comparison of FLOPs and Attention Map Ratio.
We also investigate the FLOPs and Attention Map Ratio (indicating the proportion of ’1’s in the lower triangle of the attention mask) required by the different architectures when trained on the same input data. As shown in Table 3, We find that SepLLM can significantly reduce FLOPs by approximately 30%. After plotting the loss ratios between SepLLM and Vanilla under the same FLOPs (see Figure 5(b)), we observe that SepLLM has lower loss than Vanilla. This indicates that our SepLLM architecture at least has a comparable ability to extract useful information from the dataset during training as Vanilla. Besides, the detailed wall-clock time per iteration and the real-time speed up are illustrated in Appendix C and Figure 1.
@@ 4.4 Post-training
Since training from scratch is time-consuming, we also conduct post-training experiment using 93000-step Pythia-1.4B-deduped checkpoint officially released by Pythia (see Section 4.1 for details). Figure 6 displays the loss curves for post-training, where SepLLM (n=64, larger lr) denotes that we employ an entire cosine learning rate scheduler (including a warm-up process starting from 0) identical to that of original Pythia-1.4B-deduped from step 0. SepLLM (n=64) and SepLLM (n=128) utilize a cosine learning rate scheduler that continues to decay from the 93000th step. From Figure 6, it is evident that increasing n and appropriately raising the learning rate both facilitate the decrease in loss. Moreover, this also illustrates that SepLLM for a swift transformation from a full-attention Transformer checkpoint to a model that aligns with the requirements of the SepLLM architecture’s embedding distribution through post-training.
Figure 6: Training loss curves for the post-training.
@@ 4.5 Streaming Applications
SepLLM can also adapt well to streaming applications, where infinite-length interactions may occur. Here, we follow StreamingLLM (Xiao et al., 2024b) to validate scenarios of infinite-length interactions using our Tailored Streaming Design on the commonly used PG19 dataset (Rae et al., 2020), which comprises 100 extensive literary works. The results are shown in Table 4. We can observe that for the same KV cache capacity c, the perplexity of predicting the next token through SepLLM remains consistently lower than that of streamingLLM (Xiao et al., 2024b) within the range from 1M to 4M. This once again verifies the ability of KV corresponding to separators to compress segment information and their impact on predicting the probability distribution of the next token.
@@
Table 4: The perplexity comparison on the PG19 test set (Rae et al., 2020). For fair evaluation, we keep the KV cache capacity c as 324 and keep Sink Cache a as 4 for both StreamingLLM and SepLLM.
We also test the end-to-end inference time of Vanilla, StreamingLLM and our SepLLM on PG19 (Rae et al., 2020) test set based on LlaMA-3-8B (Dubey et al., 2024). Based on the aforementioned settings, we used these LLMs to generate 20K and 64K tokens to evaluate their total inference time (wall-clock time), average perplexity, and average KV cache usage. For both SepLLM and StreamingLLM, the maximum whole KV cache capacity was set to 800 (i.e., c=800), and the initial KV size was set to 4 (i.e., a=4). For SepLLM, we additionally set s=64 and w=256. The results are shown in Table 5.
@@
Table 5: The perplexity and runing time comparison on the PG19 test set (Rae et al., 2020). r.KV means the average runtime KV cache usage in the generation process.
Above results demonstrate that our SepLLM can achieve lower perplexity with less wall-clock time as well as lower average runtime KV usage, especially for longer sequences, given the same max KV cache capacity c.
@@ 4.6 Ablation Study
We conduct various ablation experiments specifically for long streaming applications. This includes a detailed study of the impact of various hyperparameters across different text lengths (5K to 20K). The experimental results about s and (w,c) pair are illustrated in Table 6 and Table 7 respectively. The conclusions are:
@@
Table 6: The perplexity and average runtime KV cache usage of SepLLM with respect to different Separator Cache sizes (s) on WikiText (Merity et al., 2017), in which a=4, w=224, c=324.
@@
  • c and w: As can be seen in Table 7, c and w can impact the perplexity in the scenario of long streaming input with lengthy text. Moreover, as they increase, the perplexity decreases accordingly.
@@
  • Initial Tokens: Initial Tokens are crucial for modeling context in long streaming inputs, whether for SepLLM or StreamingLLM. Removing them has a significant impact on the perplexity of long texts. This conclusion is consistent with the paper (Xiao et al., 2024b).
@@ 4.7 Generalization and Information Retrieval
To verify the generalization of our SepLLM, we adapt our method to models with different architectures and different scales. The results in Appendix D can validate the generalization of our proposed framework. Specifically, we adapt our proposed SepLLM to different models including Pythia-6.9B, Pythia-12B (Biderman et al., 2023), Llama-3-8B-Base/Instruct (Dubey et al., 2024) and Falcon-40B (Almazrouei et al., 2023). Moreover, we also conduct the Needle In A Haystack experiment, which further demonstrates the compression capability of separator tokens for segment information. As illustrated in Appendix E, SepLLM can retrieve the needle in most scenarios. In comparison, StreamingLLM (Xiao et al., 2024b) cannot complete this task. Appendix F discusses the effect of separators.

### v2

# SepLLM: Accelerate Large Language Models by Compressing One Segment into One Separator

@@ 1 Introduction
  • We have made our implementation publicly available at sepllm.github.io. Our codebase supports efficient multi-node distributed training with accelerated attention module Sep-Attention and also supports numerous existing Fusion Operators to accelerate the training process, such as fused rope (Su et al., 2023), fused layer norm, etc.
@@ 3.3 Tailored Streaming Design
Figure 4 illustrates the SepLLM’s processing architecture for streaming inference applications. The diagram depicts multiple iterations, with each row representing a distinct processing step. The system simultaneously maintains four specialized cache blocks: Initial Cache, Separator Cache, Past Window Cache, and Local Window Cache. Specifically, Initial Cache captures the attention sink proposed by Xiao et al. (2024b). Local Window and Past Window Caches store the KVs for consecutive tokens, with Past Window Cache serving as an overflow buffer for the Local Window Cache. Separator Cache retains the KVs for separators which contain condensed segment information.
[…]
  • c: The maximum capacity of the entire KV cache.
@@
  • w: The maximum capacity of Local Window Cache. Notably, w is also the minimum value of n after runtime KV cache usage $Size_{\text{run}}$ reaches c for the first time.
@@
$\lim\limits_{m\to\infty}\!\overline{{Size}_{\text{run}}}$ $=\lim\limits_{m\to\infty}\!\bar{\textit{{n}}}_{m}+\textbf{{a}}+\textbf{{s}}$
@@
Our positional encoding strategy for streaming settings is the same as the state-of-the-art StreamingLLM (Xiao et al., 2024b), designed specifically for infinite-length inputs, where we focus on positions within the cache instead of those in the original text.
@@ 4.1 Experimental Settings
We evaluate our proposed SepLLM on the following tasks, i.e., training-free, training-from-scratch, post-training, and streaming applications.
@@
Two popular model families, i.e., Pythia (Biderman et al., 2023) and Llama-3 (Dubey et al., 2024), are employed for evaluation. Specifically, Pythia-160m-deduped is used as the backbone in the training-from-scratch tasks since the model, data, configurations, and checkpoints are all open-source and the training results are reproducible. As for post-training settings, we take Pythia-1.4B-deduped as our backbone model. Even though Llama-3 exhibits powerful performance on various downstream tasks, the training experimental details are not available. Therefore, we only use Llama-3 for training-free and streaming tasks.
@@
In the training-from-scratch and post-training tasks, the deduplicated Pile (Gao et al., 2020) is utilized for training, which contains about 207B tokens. And all other configurations are the same as the corresponding settings as Pythia (Biderman et al., 2023). Specifically, the training epoch is set to 1.5 epoch (143000 steps with the global batch size as 1024), which means about 300B tokens in total are utilized for training from scratch, which is identical to Pythia (Biderman et al., 2023).
@@
Table 1: Evaluation results and average runtime KV cache usage for training-free experiments on GSM8K-CoT 8-shots and MMLU 5-shots. For SepLLM and StreamingLLM, three initial tokens’ KV is kept for this experiment. r.KV(%) here represents the ratio of KV usage at runtime for the respective method compared to Vanilla.
@@ 4.2 Training-free
The representative and commonly-used GSM8K-CoT (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021)) are adopted. GSM8K-CoT (Cobbe et al., 2021) tests a model’s ability to solve mathematical problems by evaluating its reasoning and step-by-step problem-solving skills. CoT (Wei et al., 2022) means the ability to simulate a reasoning process by breaking down complex problems into a series of logical steps. And the default 8 shots are adopted. MMLU (Hendrycks et al., 2021) assesses a model’s general knowledge and reasoning ability across a wide range of subjects, such as history, science, mathematics and so on. The commonly-used 5-shot setting is used for MMLU.
@@
Table 2: The performance of downstream tasks and the average runtime KV cache usage in the training-from-scratch setting.
@@ Results.
The experimental results for training-free are shown in Table 1. “Vanilla” represents the original Llama-3 model with full attention, while “StrmLLM” represents StreamingLLM (Xiao et al., 2024b). n means the number of KV for Neighboring Tokens we retain. For SepLLM, all the KV for Separator Tokens are kept and for the setting SepLLM (n=256), we find that SepLLM exhibits comparable performance in both multi-step mathematical CoT task and multidisciplinary knowledge reasoning tasks, when compared to the full-attention Llama-3. SepLLM achieves this using only 47.36% of the KV utilized by the original Llama-3 for reasoning, indicating SepLLM’s capability of modeling both the contexts requiring multi-step logical analysis and those involving multi-domain knowledge reasoning while retaining only 50% original KV.
StrmLLM (n=256) setting corresponds to removing all separators’ KV from  SepLLM (n=256) setting, except for those in Neighboring and Initial tokens. We observe a noticeable decrease in both mathematical analysis and multidisciplinary knowledge reasoning abilities for StrmLLM (n=256). StrmLLM (n=256) utilizes only 26.00% and 37.73% of the KV for the GSM8K and MMLU tasks, respectively, which are less than  SepLLM (n=256) (47.36% and 44.61% respectively). Consequently, we increase the n of StrmLLM to 380, aligning the kept KV on GSM8K to be equal to  SepLLM (n=256) (approximately 47%, while on MMLU task, StrmLLM (n=380) retains 52.5% of the KV, significantly higher than  SepLLM (n=256)). This leads to improved performance compared to StrmLLM (n=256). However, it still remains lower than the full-attention Llama-3 and  SepLLM (n=256). This indicates that the KV of separators indeed encapsulates information contained within their respective segments, and removing them significantly impacts the Transformer’s understanding and reasoning abilities.
@@
Figure 5: Training loss curves for training from scratch. 5(b) shows the ratios of the loss values of different methods to that of Vanilla with respect to FLOPs.
@@ 4.3 Training from Scratch
We train the original Pythia-160m-deduped model as well as the Pythia-160m-deduped model modified with the SepLLM (and StreamingLLM) architecture on the Pile dataset for 143,000 steps using a global batch size of 1024 (involving approximately 300B tokens in total for training). All training configurations are consistent with Pythia (Biderman et al., 2023) (see Section 4.1). And following Pythia (Biderman et al., 2023), we conduct tests on the following downstream tasks: ARC-Challenge and ARC-Easy (Clark et al., 2018), LAMBADA (Paperno et al., 2016) (for Perplexity and Accuracy), LogiQA (Liu et al., 2021), PIQA (Bisk et al., 2020), SciQA (Welbl et al., 2017). From the loss curves depicted in Figure 5 and the downstream performance in Table 2, we draw the following analysis.
@@
We find that employing a certain hybrid architecture is beneficial to both the training loss and the performance on downstream tasks. For instance, by modifying only the first self-attention layer to full attention in the experiment corresponding to SepLLM (n=64) (denoted as SepLLM (n=64,H)), there is a moderate optimization in both the training process and downstream tasks. If both the first and last attention layers are changed to full attention (denoted as SepLLM (n=64,H/T)), this optimization becomes more pronounced. For example, LAMBADA perplexity decreases from 40.08 for SepLLM (n=64) to 36.54 for SepLLM (n=64,H) and 33.41 for SepLLM (n=64,H/T).
@@
The experiment with the setting StrmLLM (n=64) corresponds to SepLLM (n=64), but does not consider separators other than those in Neighboring and Initial tokens. We observe a significant slowdown in the training loss decrease for StrmLLM (n=64), and its performance deteriorates across various downstream tasks. This indicates that the KV corresponding to separators indeed contain information about the segments they belong to, which are beneficial to predicting subsequent tokens.
@@
Table 3: The comparison of FLOPs and Attention Map Ratios.
We also investigate the FLOPs and Attention Map Ratio (indicating the proportion of ’1’s in the lower triangle of the attention mask) required by the different architectures when trained on the same input data. As shown in Table 3, We find that SepLLM can significantly reduce FLOPs by approximately 30%. After plotting the loss ratios between SepLLM and Vanilla under the same FLOPs (see Figure 5(b)), we observe that SepLLM has lower loss than Vanilla. This indicates that our SepLLM architecture at least has a comparable ability to extract useful information from the dataset during training as Vanilla. Besides, the detailed wall-clock time per iteration and the wall-clock time speedups are illustrated in Appendix C and Figure 1.
@@ 4.4 Post-training
Since training from scratch is time-consuming, we also conduct post-training experiments using 93000-step Pythia-1.4B-deduped checkpoint officially released by Pythia (see Section 4.1 for details). Figure 6 displays the loss curves for post-training, where SepLLM (n=64, larger lr) denotes that we employ an entire cosine learning rate scheduler (including a warm-up process starting from 0) identical to that of original Pythia-1.4B-deduped from step 0. SepLLM (n=64) and SepLLM (n=128) utilize a cosine learning rate scheduler that continues to decay from the 93000th step. From Figure 6, it is evident that increasing n and appropriately raising the learning rate both facilitate the decrease in loss. Moreover, this also illustrates that SepLLM can achieve a swift transformation from a full-attention Transformer checkpoint to a model that aligns with the requirements of the SepLLM architecture’s embedding distribution through post-training.
Figure 6: Training loss curves for the post-training setting.
@@ 4.5 Streaming Applications
SepLLM can also adapt well to streaming applications, where infinite-length interactions may occur. Here, we follow StreamingLLM (Xiao et al., 2024b) to validate the scenarios of infinite-length interactions using our Tailored Streaming Design on the commonly used PG19 dataset (Rae et al., 2020), which comprises 100 extensive literary works. The results are shown in Table 4. We can observe that for the same KV cache capacity c, the average perplexity of predicting the next token through SepLLM remains consistently lower than that of streamingLLM (Xiao et al., 2024b) within the range from 1M to 4M input length. This once again verifies the ability of KV corresponding to separators to compress segment information and their impact on predicting the probability distribution of the next token.
@@
Table 4: The perplexity comparison on the PG19 test set (Rae et al., 2020). For fair evaluation, we keep the entire KV cache capacity c as 324 and Initial Cache capacity a as 4 for both StreamingLLM and SepLLM. w=224, s=32/64 for SepLLM.
We also test the end-to-end inference time of Vanilla, StreamingLLM and our SepLLM on PG19 (Rae et al., 2020) test set based on LlaMA-3-8B (Dubey et al., 2024). Based on the aforementioned settings, we used these LLMs to generate 20K and 64K tokens to evaluate their total inference time (wall-clock time), average perplexity, and average runtime KV cache usage. For both SepLLM and StreamingLLM, the maximum whole KV cache capacity was set to 800 (i.e., c=800), and the Initial Cache capacity was set to 4 (i.e., a=4). For SepLLM, we additionally set s=64 and w=256. The results are shown in Table 5.
@@
Table 5: The average perplexity and running time comparison on the PG19 test set (Rae et al., 2020). r.KV means the average runtime KV cache usage in the generation process.
Those results demonstrate that our SepLLM can achieve lower perplexity with less wall-clock time as well as lower average runtime KV usage, especially for longer sequences, given the same max KV cache capacity c.
@@ 4.6 Ablation Study
We conduct various ablation experiments specifically for long-input applications. This includes a detailed study of the impact of various hyperparameters across different text lengths (5K to 20K). The experimental results about s and (w,c) pair are illustrated in Table 6 and Table 7 respectively. The conclusions are:
@@
Table 6: The perplexity and average runtime KV cache usage of SepLLM with respect to different Separator Cache capacities (s) on WikiText (Merity et al., 2017), in which a=4, w=224, c=324.
@@
  • c and w: As can be seen in Table 7, c and w can impact the average perplexity in the scenario of long streaming input with lengthy text. Moreover, as they increase, the perplexity decreases accordingly.
@@
  • Initial Tokens: Initial Tokens are crucial for modeling context of long streaming inputs, whether for SepLLM or StreamingLLM. Removing them has a significant impact on the perplexity of long texts. This conclusion is consistent with the paper (Xiao et al., 2024b).
@@ 4.7 Generalization and Information Retrieval
To verify the generalization of our SepLLM, we adapt our method to models with different architectures and different scales. The results in Appendix D can validate the generalization of our proposed framework. Specifically, we adapt our proposed SepLLM to different backbones including Pythia-6.9B, Pythia-12B (Biderman et al., 2023), Llama-3-8B-Base/Instruct (Dubey et al., 2024) and Falcon-40B (Almazrouei et al., 2023). Moreover, we also conduct the Needle In A Haystack experiment, which further demonstrates the compression capability of separator tokens for segment information. As illustrated in Appendix E, SepLLM can retrieve the needle in most scenarios. In comparison, StreamingLLM (Xiao et al., 2024b) cannot complete this task. Appendix F discusses the effect of separators.